The base of a triangle is 2 MORE than TWICE the height of the triangle. Find the measure of the base and the height if the area is 20ft2 (20 feet square).
Answer by edjones(8007)

b=2h+2
.5bh=A
.5(2h+2)h=20
.5(2h^2+2h)=20
h^2+h=20
h^2+h-20=0
(h+5)(h-4)=0
h=4, b=10
.
Ed
.
Now a question for you.
What's the difference between "...square feet" and "...feet square"
This is a very important distinction.